QUESTION 10 The following trial balance has been extracted from the books of SEREBOUR ENT. on 30 November 2003 Sales Returns Debtors and creditors Office equipment 125,658 1,902 7,983 6,341 11,257 - cost 10,000 accumulated depreciation 11/2002 1,550 Vehicle -cost 3,500 accumulated depreciation 1/1/2002 700 Purchases Stock at 1/1/2002 Carriage inwards Carriage outwards Vehicle expenses Electricity Wages and salaries Rent and business rates Stationery and postages Bank deposit account Bank Discount allowed and received VAT creditor 64,726 5,000 908 272 1,349 1,803 11,550 8.800 2,681 10,000 4,797 5,652 3,765 1,325 PAYE creditor Capital at 1 December 2002 Drawings 453 4,300 15,000 oan 16.000 163,636 163.636 Following the extraction of the above trial balance from the ledgers and an examination of the accounting records, the following additional information were obtained.

PAYE creditor Capital at 1 December 2002 Drawings Loan 453 4,300 15,000 16,000 163,636 163.636 Following the extraction of the above trial balance from the ledgers and an examination of the accounting records, the following additional information were obtained. 1. The value of the stock on hand at 30 November 2003 was c5,700 2. The bank deposit was made on 1 June 2003 . This account earns interest at 8% per annum. The balance shown in the ledgers is the only deposit into this account. Interest is credited on 31 December annually 3. Bank charges accrued on 30 November 2003 are estimated to be 60 4. Vehicle expenses include road tax of c125 per annum for a vehicle, which expires on 31 December 2003, and vehicle insurance of c360 per annum which was paid on 1 July 2003 5. There is one employee whose gross wage is c1,050 per month. The wage for November 2003 has not yet been entered in the ledgers. 6. The balance of the rent and rates account includ 31 December 2003 of c1,200 and payment of rates for the six months to 31 March 2003 of 4,000 es a payment of rent for the three months to 7. Following a review of debtors at the year end, it is decided that a bad debt of c1,207 should be written off, and a provision for doubtful debts made of 2% of debtors 8. Office equipment is to be depreciated at 10% per annum on cost, and motor vehicles are to be depreciated at 20% per annum on cost. 9. Loan interest of 10% per annum is to be accrued. Requirements Using the above information, prepare a trading and profit and loss account for SEREBOUR ENTERPRISE for the period ended 30 November 2003 together with the balance sheet at that date. (20 marks)
Answer #1

SEREBOUR ENTERPRISE

Trading & Profit & Loss Account for the year ending 30th Nov 2003

Particulars

Amount

Particulars

Amount

Opening stock

Purchases      64726

Less: Returns 6341

Carriage inwards

Wages & Salaries   11550

Add: Outstanding 1050

Gross Profit C/d

Interest on loan

Carriage outwards

Electricity

Stationery & postages

Discount allowed

Bank charges

Bad debts Written off

Provision for doubtful debts

Depreciation :

Office equipment     917

Motor vehicles           642

------------------

Vehicle Expenses          1349

Less : prepaid road tax 10

         : prepaid insurance 210

-------------------------

Rent & business rates     8800

Less : prepaid                     400

Less: Outstanding paid     2000

--------------------

            Net profit c/d

5000

58385

908

12600

52563

-------------

129456

=========

1467

272

1803

2681

5652

60

1207

201

1559

1129

6400

34364

---------

56795

---------

Sales      125658

Less: Returns 1902

Closing stock

Gross Profit b/d

Interest on deposit

Discount Received

         

  

Net profit b/d

123756

5700

------------

129456

=========

52563

467

3765

56795

---------

34364

Calculation of Provision for doubtful debts:

11257-1207= 10050

10050*2/100=201

Calculation of Depreciation:

Office equipment = 10000*10/100=1000*11/12=917

Motor vehicles       =3500*20/100=700*11/12=642

Balancesheet as on 30th Nov 2003

Liabilities

Amount

Assets

Amount

Capital                   4300

Less : Drawings   15000

Add: Net profit   34364

                              -------------

Loan     16000

Add : interest on loan 1467

Creditors

Wages outstanding

VAT Creditor

PAYE Creditor

23664

17467

7983

1050

1325

453

Vehicle 3500

Less: depreciation 1342

Office Equipment     10000

Less: depreciation 2467

Bank deposit 10000

Add: interest 467

Bank

Debtors      11257

Less : Bad debts 1207

Less : provision 201

Closing stock

Prepaid road tax

Prepaid insurance

Prepaid rent

2158

7533

10467

4797

9849

5700

10

210

400

NB: Data inappropriate

Financial Year(period) is not given

Thankyou